OpenStudy (anonymous):

Solve for V in V = s3, if s = 4.

OpenStudy (jackellyn):

is that s^3? or s*3? V=s*3 V= (4)3 V =12 V = s^3 V= (4)^3 V = 64

OpenStudy (anonymous):

i bet it is the second one

OpenStudy (turingtest):

@elitepenguin if you mean\[V=s^3\]please type it as V=s^3 in your post to avoid confusion with multiplication thanks